# Survey on Large Language Model-Enhanced Reinforcement Learning: Concept, Taxonomy, and Methods

### v1

@@ VIII-A Applications of LLM-Enhanced RL
  • Autonomous Driving: Autonomous driving uses RL to make decisions in complex, ever-changing environments that involve understanding both sensor data (visual, lidar, radar) and contextual information (traffic laws, human behavior). LLM-enhanced RL could employ LLMs to 1) process this multi-modal information and natural language instructions; or 2) design comprehensive rewards based on multi-disciplinary metrics such as safety, efficiency, and passenger comfort.
@@
VIII-B Opportunities for LLM-Enhanced RL
@@
  • RL: Existing work such as [112][108][107] mainly focus on the general RL while various specialized branches of RL are still under-exploited, e.g. multi-agent RL, safe RL, transfer RL, explainable RL, and other niches within the field. In multi-agent RL, LLM-enhanced strategies could be employed to facilitate communication and collaboration among RL agents. The natural language understanding capabilities of LLMs can be used to interpret and generate instructions or strategies among agents, enhancing cooperative behaviors or competition strategies; Safe RL could benefit from the reasoning and predictive capabilities of LLMs to design cost functions that encourage safety criteria compliance, reducing the risks of dangerous exploration; In transfer RL, LLMs can assist in identifying similarities between tasks or environments, leveraging their vast knowledge base to facilitate knowledge transfer and thus improve learning efficiency and adaptability across different tasks.
  • LLM: Present methodologies such as [102][39][90] predominantly harness LLMs by employing straightforward prompting techniques to elicit responses. However, integration of LLMs with the tool manipulation capability unlocks new possibilities of LLMs. Tools such as mathematical solver and internet browser, retrieval-augmented generation (RAG) techniques, and API assistance can be leveraged to equip LLMs with a plethora of external knowledge and computational capabilities, improving the performance of LLM-enhanced RL in tasks that require real-time data fetching, external knowledge and mathematical model solving.
VIII-C Challenges of LLM-Enhanced RL
@@
While LLM improves different issues in the RL paradigm, the success of LLM-enhanced RL is inherently tied to the capabilities and limitations of the underlying LLM. The primary concerns revolve around the dependency on LLM capabilities, interaction efficiency, and the broader implications of deploying such systems in real-world scenarios.
  • Dependence on LLM’s Capabilities: The LLM-enhanced RL paradigm depends on the capabilities of LLMs, which affects the performance of RL agent. Biases and hallucination inherent in pre-trained LLMs can lead to skewed or inaccurate interpretations of multi-modal inputs. Additionally, despite the vast knowledge base of LLMs, they may struggle to adapt to highly specific or novel task environments not well-represented in their training data. Therefore, to effectively integrate LLMs into RL paradigm, a comprehensive evaluation is required in advance.
  • Interaction Efficiency: The integration of LLMs into the RL learning process introduces complexities in terms of computational overhead and response times, which slow down the learning process of RL. These complexities require advances in both algorithms and hardware implementation for faster inference with fewer computational resources.
  • Ethical, Legal, and Safety Concerns: In practical usages, the use of LLMs involves complex ethical, legal, and safety concerns. Data privacy, intellectual property, and accountability for AI decisions should be carefully discussed. In addition, LLMs in RL introduce new sectors for adversarial attacks, potentially compromising the integrity and safety of the system.
@@ IX Conclusion
  • Information Processor: LLMs extract observational representations and formal specification languages for RL agents, thus increasing the sample efficiency. Future directions include incorporating goal-based information and integrating multi-modal environmental information to obtain stronger information extraction ability.

### v3

@@ VIII-B Applications of LLM-Enhanced RL
  • Autonomous Driving: Autonomous driving uses RL to make decisions in complex, ever-changing environments that involve understanding both sensor data (visual, lidar, radar) and contextual information (traffic laws, human behavior). LLM-enhanced RL could employ LLMs to 1) process this multimodal information and natural language instructions; or 2) design comprehensive rewards based on multi-disciplinary metrics such as safety, efficiency, and passenger comfort.
@@
VIII-C Opportunities for LLM-Enhanced RL
@@
  • RL: Existing work such as [138, 134, 129] mainly focus on the general RL while various specialized branches of RL are still under-exploited, e.g. multi-agent RL, safe RL, transfer RL, explainable RL, multi-task RL, in-context RL and human-centric RL. In the multi-agent area, compared to various works about multi-LLM-agent collaboration [165][166], LLM-based multi-agent RL remains largely unexplored [167]. A recent survey discussed some open research problems within this field [168]. LLM-enhanced strategies could be employed to facilitate communication and collaboration among RL agents. The natural language understanding capabilities of LLMs can be used to interpret and generate instructions or strategies among agents, enhancing cooperative behaviors or competition strategies; Safe RL could benefit from the reasoning and predictive capabilities of LLMs to design cost functions that encourage safety criteria compliance, reducing the risks of dangerous exploration; In transfer RL, LLMs can assist in identifying similarities between tasks or environments, leveraging their vast knowledge base to facilitate knowledge transfer and thus improve learning efficiency and adaptability across different tasks; For multi-task RL, LLMs enhance agents by processing diverse entity representations and aligning instructions (Information Processor), generating reduced action sets and expert actions for various tasks (Decision-Maker), and creating task-specific trajectories and dynamic representations (Generator). Recently, in-context RL has emerged as a promising field by leveraging the in-context learning capability of LLMs to solve decision-making problems [169]. Given a query state and an in-context offline dataset, a trained LLM exhibits both online exploration and offline conservation while avoiding extensive trial and error and is sample-efficient. Additionally, human-centric RL is another prominent field in healthcare and robotics, where AI and humans learn and communicate with each other for collaboration [170]. In this setting, an LLM naturally serves as a perfect mediator to facilitate bidirectional communication through natural language interactions.
  • LLM: While LLMs have been integrated with RL in various methods, several promising directions remain to be explored to further enhance LLM capabilities for RL. It can be divided into the language model view and the language agent view as follows:
    • From the model perspective: LLM could be enhanced by an external knowledge base and continual learning. Retrieval-augmented generation (RAG) techniques help LLM to retrieve the most relevant data in an external database, which could be used to attach knowledge in the task domains [171]. Continual learning techniques are also a hot field that enables models to continuously acquire new knowledge while retaining previously learned capabilities [172]. This technique allows both LLMs and RL agents to continuously evolve and adapt to new tasks or environments while maintaining their fundamental pre-trained abilities, leading to more flexible and robust learning systems.
    • From the agent perspective: Equipping LLM-based agents with specialized modules—namely, planning modules, memory modules, and action modules—can significantly enhance the capabilities of LLMs [173, 174]. In the planning module, multi-step reasoning techniques such as CoT-based reasoning [74] and Monte Carlo tree search (MCTS) [175] can be used to enhance LLM’s long-term planning ability, improving the long-term task decomposition ability of RL agents; for the memory module, long- and short-term human-memories and corresponding memory retrieval mechanisms provide a way to store previous experiences and learn adaptively along with the RL agents [176]; for the action module, tool integration presents another promising direction to unlock new possibilities for LLMs. External tools such as mathematical solvers [177] and internet browsers [178] can augment LLMs’ capabilities in RL tasks requiring complex mathematical computations and real-time information processing, respectively. Furthermore, collaboration of multiple LLM agents [179] is another promising direction. With each LLM playing different roles and acquiring different information, multiple distinctive LLM agents solve complex problems by communicating and exchanging information. In the context of LLM-enhanced RL, different LLMs could potential serves different roles and work together to guiding RL. For example, in a group of three LLM agents, one LLM focuses on guiding short-term or immediate problem-solving for RL; another LLM is responsible for long-term planning for RL; the last LLM serves as a coordinator responsible for overseeing the group and adjusting horizon consideration accordingly.
VIII-D Challenges of LLM-Enhanced RL
@@
While LLM improves different issues in the RL paradigm, the success of LLM-enhanced RL is inherently tied to the capabilities and limitations of the underlying LLM. The primary concerns revolve around the inherent limitations of LLMs, adaptability of LLMs in RL environments, computational demands, and the broader implications of deploying such systems in real-world scenarios.
  • Inherent Limitations of LLMs: The effectiveness of LLM-enhanced RL systems is fundamentally constrained by the inherent capabilities and limitations of the underlying language models. The presence of systematic biases and potential hallucinations in pre-trained LLMs can significantly impact the reliability of multimodal input interpretation, potentially compromising the overall performance of the RL agent. This necessitates the development of robust evaluation frameworks to systematically characterize and delineate the capability boundaries of LLMs given specified RL contexts. Furthermore, incorporating uncertainty quantification methods aids in identifying unreliable answers, increasing the trustworthiness of LLMs’ responses [180].
  • Adaptability of LLMs in RL Environments: Despite the vast knowledge base of LLMs, they may struggle to adapt to specific or novel RL task environments that are not well-represented in their training data. This calls for methods to expand the task-related knowledge for LLM and ground LLMs’ inherent knowledge in specific domains. In terms of expanding the task-related knowledge, RAG-related techniques attach LLMs with the external domain-knowledge without further training [171]. When fine-tuning LLMs, employing data augmentation techniques such as generating synthetic data is another way to expand the diversity of training scenarios and improve generalization to novel environments [181]. Additionally, continual learning mechanisms help preserve previously acquired knowledge while adapting to new information [182]. For domain-specific knowledge grounding, approaches include training value functions to evaluate the long-term utility of LLM instructions [83]. Expert trajectories provide another valuable source of information, allowing LLMs to learn optimal decision-making patterns through in-context learning.
  • Computational Demands: The integration of LLMs into the RL learning process introduces complexities in terms of computational overhead and inference times, which slow down the learning process of RL. To address this challenge, several potential solutions have been proposed across different levels. At the data level, input compression techniques such as prompt pruning [183] can be employed to directly shorten the model input, significantly reducing inference time without substantial loss in performance. At the model level, efficient architecture designs like mixture-of-experts (MoE) [184] enable conditional computation, activating only relevant parts of the model for each input, thus reducing computational costs. Similarly, structured state space models (SSM) [185] offer linear-time complexity for sequence modeling, providing a more efficient alternative to traditional attention mechanisms. At the system level, advanced caching strategies [186] and asynchronous processing techniques [187] can be implemented to reuse intermediate computations and parallelize the execution, respectively.
  • Ethical, Legal, and Safety Concerns: In practical usages, the use of LLMs involves complex ethical, legal, and safety concerns. Data privacy, intellectual property, and accountability for AI decisions should be carefully discussed. To address these issues, researchers are developing robust frameworks for responsible AI deployment. At the privacy level, differential privacy techniques [188] are being implemented to protect individual data during training and inference. For transparency, efforts focus on developing interpretable AI systems, allowing stakeholders to audit AI-driven decisions [189]. To enhance system robustness, adversarial training methods are being explored to strengthen LLM-RL systems against potential attacks [190]. Regarding ethical and legal issues, a recent survey also analyzed potential solutions to integrate ethical standards and societal values into LLM  [191].
@@ IX Conclusion
  • Information Processor: LLMs extract observational representations and formal specification languages for RL agents, thus increasing the sample efficiency. Future directions include incorporating goal-based information and integrating multimodal environmental information to obtain stronger information extraction ability.